Primary tabs

Story Archive

Category Title Date Excerpt Attached Images
Current Events Forty Years of Beam! 17 December 2014 It was 40 years ago that the laboratory celebrated extraction of first beam from the 500 MeV cyclotron, the same cyclotron driving science at the lab today. The TRIUMF Alumni & Retirees Association visited to connect and learn of the present status of the lab but also to share memories of times past.
Current Events Beam delivery update: Last update of 2014? 12 December 2014 The cyclotron is still delivering close to 100 µA to BL2C4 for isotope production and protons as needed to BL1B for PIF. BL2A remains off while the next experiment is being set up.
Current Events December 12 ARIEL Snapshot 12 December 2014 Productive e-linac lessons-to-learn session held on Dec. 10th. High-level application successfully applied to beam experiments, accepted by XAL Collaboration for public release. Held review of EHBT optics in the BL4N era. Ongoing work on the ARIEL Target pit layout and planning of the target station. Release of ARIEL driver beam and target specifications. 
Current Events Beam delivery update 11 December 2014 They cyclotron’s running fairly well, delivering close to 100 µA to BL2C4 for isotope production and protons as needed to BL1B for PIF. BL2A is off at least overnight as we have no RIB users at ISAC.
Current Events Beam delivery update 10 December 2014 The cyclotron is delivering 9.5 µA to BL2A, ~95 µA to BL2C4 for isotope production, and beam as needed to BL1B for PIF. BL2A may go off for a few hours if RIB is not needed overnight, but otherwise we expect to continue like this for the next few days.
Current Events Beam delivery update 09 December 2014 Today’s planned beam development is still underway but is scheduled to wrap up by 20:30.
Current Events Beam delivery update: Short maintenance and beam development tomorrow 08 December 2014 The cyclotron is currently delivering ~110 µA to BL1A, 9.8 µA to BL2A, and nA as needed to BL2C1 for PIF. This will continue until ~07:00 tomorrow, at which point the cyclotron will go off for a short maintenance period followed by beam development.
Current Events Beam delivery update 05 December 2014 The BL1A current has been reduced to ~110 µA. BL2A is running at 9.8 µA while BL2C1 continues to take nA as needed for PIF. This is all expected to continue through the weekend.
Current Events Canada Launches ST&I Strategy 05 December 2014 Prime Minister Stephen Harper and Minister of State (S&T) Ed Holder announced the release of an updated federal science, technology and innovation strategy, with a clear vision for Canadian excellence.
Current Events December 5 ARIEL Snapshot 05 December 2014  The e-gun solenoid faulty field clamps were replaced and solenoid was re-installed. Beam dump modified lower center block and insert installed, ready for leak check. Planning for cable tray system for 284' mezzanine, North and South side of ARIEL tunnel.

Pages

Subscribe to Recent Stories